OpenStudy (volleyballlover55):

Which expression is equivalent to 5^4 ⋅ 5^–6? 5^–10 5^6 fraction 1 over 5 to the power 2 fraction 1 over 5 to the power negative10

OpenStudy (volleyballlover55):

@satellite73

OpenStudy (volleyballlover55):

25

OpenStudy (anonymous):

oh wait i read it wrong, sorry, skip that

OpenStudy (volleyballlover55):

ok

OpenStudy (anonymous):

add the exponents is all \[\huge 5^5\times 5^{-6}=5^{5+(-6)}\]

OpenStudy (volleyballlover55):

5^-11? @satellite73

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

\[5-6\neq -11\]!

OpenStudy (volleyballlover55):

oh so its -1

OpenStudy (anonymous):

aye yikes it is actually \[5^4\times 5^{-6}\]which is \[5^{4-6}=5^{-2}\]

OpenStudy (anonymous):

better known as \[\frac{1}{5^2}\]

OpenStudy (volleyballlover55):

ok!